OpenStudy (anonymous):

Perform the operation and simplify the result if possible. (2 m ^2 - 9) / (m^4 - 25) + (4 - m^2) / (m^4 - 25)

OpenStudy (freckles):

add the fractions (you have the same denominator) then once you have added factor numerator and denominator to see if their are any common factors

OpenStudy (anonymous):

I got m^2 -5/m^4 -25

OpenStudy (freckles):

\[\frac{m^2-5}{m^4-25}\] great factor the denominator you should see a common factor in the top and bottom after doing this

OpenStudy (freckles):

think difference of squares

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

\[\frac{ 2m ^{2} }{ (m ^{2}+5)(m ^{2}-5) }-\frac{ 9 }{ (m ^{2}+5)(m ^{2}-5) }+\frac{ 4 }{ (m ^{2}+5)(m ^{2}-5) }-\frac{ m ^{2} }{ (m ^{2}+5)(m ^{2}-5) }\] This is how it starts out

OpenStudy (freckles):

ok so @Austin1617 as shown you how to factor the denominator you should get \[\frac{m^2-5}{(m^2-5)(m^2+5)} \] you should see something that cancels @ludvic and what mean by cancel here is that something divided by the same something is 0 provided that something isn't 0 itself for example we know 4/4=1 -3/-3=1 and so on... a/a=1 for all a excluding a=0 so this is what we mean by canceling here \[\frac{m^2-5}{(m^2-5)(m^2+5)}=\frac{m^2-5}{m^2-5} \cdot \frac{1}{m^2+5}=...\]

OpenStudy (anonymous):

it's m^2 +5 as the answer ? @freckles

OpenStudy (freckles):

well why did you put m^2+5 on top? it is on the bottom just a second ago?

OpenStudy (anonymous):

so you put 1 over m^2 +5?

OpenStudy (freckles):

yep because that is what is left behind after replacing (m^2-5)/(m^2-5) with 1

OpenStudy (freckles):

and I think you mean 1 over (m^2+5)